For each line, use the given conditions to write an equation in the point-slope form and slope-intercept form.
1. Slope = 4, passing through (1,3)
2. Slope = 8, passing through (4,-1)
3. Slope = -5, p[assing through (-4,-2)
4. Slope = -2, passing through (0,-3)

Answer by stanbon(75887)

For each line, use the given conditions to write an equation in the point-slope form and slope-intercept form (y = mx+b).
---
Solve for "b" on each of these problems:
----
1. Slope = 4, passing through (1,3)
3 = 4*1+b
b = -1
Equation: y = 4x-1
--------------------------
2. Slope = 8, passing through (4,-1)
-1 = 8*4 + b
b = -33
Equation: y = 8x-33
--------------------------
3. Slope = -5, p[assing through (-4,-2)
-2 = -5*-4 + b
b = -22
Equation: y = -5x-22
=========================
4. Slope = -2, passing through (0,-3)
-3 = -2*0 + b
b = -3
Equation: y = -2x -3
=========================
Cheers,
Stan H.
===========